### v1

# A Fast Algorithms for Fourier extension based on boundary interval data

Abstract
This paper present a new algorithm for the computation of Fourier extension based on boundary data, which can obtain a super-algebraic convergent Fourier approximation for non-periodic functions. The algorithm calculates the extension part through boundary data and connects it with the original function to form a periodic smooth function. By testing the key parameters involved, their impact on the algorithm is clarified and the optimization setting scheme of the parameters is proposed. Compared with FFT, the algorithm only needs to increase the computational complexity by a fixed small amount.
keywords:
Fourier Series, Fourier Extensions, Fourier Continuation
{AMS}
42A10, 65T40, 65T50
1 Introduction
Fourier series approximation for smooth periodic functions is popular because it can offer spectral convergence and be computed numerically via the FFT. However, the situation is changed completely once periodicity is lost. A jump discontinuity is imposed in the approximation at the domain boundaries due to the well known Gibbs phenomenon. Many ways have been developed to ameliorate the ill effects of the Gibbs phenomenon, including the methods based on Gegenbauer polynomials [9, 10, 16], the methods utilizing Padé approximations [7, 8] and the Fourier extension (or continuation) methods [4, 11, 12, 14, 13, 15, 6, 17].
The idea of the existing discrete Fourier extension can be described as follows: Let $f$ be a non-periodic, smooth and sufficiently differentiable function over $[-1,1]$, which is sampled in a predefined set of points $\{t_{\ell}\}\subset[-1,1]$, the discrete Fourier extension $C(f)$ of $f$ is defined by
$C(f)=\sum_{k=-N}^{N}{\bf c}_{k}e^{\text{i}\frac{k\pi}{T}t},\quad T>1,$ (1)
where the coefficients $\{{\bf c}_{k}\}_{k=-N}^{N}$ are chosen to minimize the error
$\min\sum_{\ell}\left|f(t_{\ell})-\sum_{k=-N}^{N}{\bf c}_{k}e^{\text{i}\frac{k\pi}{T}t_{\ell}}\right|^{2}.$ (2)
To obtain a solution to the minimization problem (2), a severely ill-conditioned least-squares system need to be solved. In [4], Boyd has introduced the truncated singular value decomposition (TSVD) to stabilize the calculation process, the accuracy almost up to machine precision. A convergence analysis by Adcock et. al states that for analytic functions the method is at least superalgebraically convergent [2]. Utilizing randomized algorithms, Lyon in [12] provided a fast algorithm for the computation with $T=2$. And in [14], Matthysen et. al presented fast algorithms for arbitrary $T$, exploiting it as a specific variant of bandlimited extrapolation. This idea was further developed in the literature [15] to form a fast algorithm for function approximation on 2D domains.
It should be noted that the goal of the extension process is to effectively connect the data at both ends of the function. We construct the following two functions:
$f_{1}(t)=\left\{\begin{aligned} &-1,&x<-\frac{1}{2}\\ &\tanh\left(\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2}\\ &1,&t>\frac{1}{2}\end{aligned}\right.$ (3)
$f_{2}(t)=\left\{\begin{aligned} &-1,&t<-\frac{1}{2}\\ &\tanh\left(100*\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2}\\ &1,&t>\frac{1}{2}\end{aligned}\right.$
Since the data of the two functions near the endpoints are the same, the same connection data can be used on the interval $[-T,1)$ and $(1,T]$. An illustration is shown in Fig. 1. This raises an issue: does the data inside the interval really help the calculation of the extension? In fact, A very effective method, known as FC-Gram method has been developed earlier, which uses only boundary data (projecting it into a basis of orthogonal polynomials) for extension calculations [5]. Recently, this method has also been extended to the approximation of two-dimensional functions[6]. The numerical implementation of the FC-gram algorithm requires a high-precision environment and it is not superalgebraic convergent.
Figure 1: Periodic extensions of $f_{1}(t)$ and $f_{2}(t)$.
This paper proposes a new fast algorithm for the calculation of Fourier extension, which has the following traits:
  • Only a few discretization points in the boundary interval are used to calculate the extension part, so the method can be computed with FFT speed. Unlike the FC-Gram algorithm, the method in this paper does not need to introduce polynomials.
  • Since the extension part only uses boundary data for calculation, the influence of the singular points inside the function on the approximation is greatly weakened. Therefore, for functions with smooth boundaries and internal oscillations, the proposed algorithm has a smaller re-solution constant than the original full-data algorithm.
All the results in this paper are based on numerical experiments, and we leave the theoretical proof as a topic for future work. The proposed algorithm has good performance and gives clear conclusions on the settings of various parameters. Therefore, we feel that it is of substantial interest for those who use FEs in practice.
1.1 Overview of the paper
In Section 2, we outline the algorithm and explain its structure. The specific calculation method of the extension part based on boundary data is given in Section 3. Section 4 systematically tests the key parameters, points out their influencing factors, and gives the optimization setting scheme. In Section 5, we give the numerical test of the final algorithm and compare it with two other fast algorithms.
2 The outline of the boundary interval algorithm
In this paper, we restrict ourselves to the case where the sampling points are uniform, that is,
$t_{\ell}=\frac{\ell}{M},\quad\ell=-M,\ldots,M.$ (4)
Let the positive integers $m_{\Delta}$ be the number of nodes taken in the boundary interval, and $S_{l},S_{r}$ be the set of corresponding subscripts,
$S_{l}=\{-M,-M+1,\ldots,-M+m_{\Delta}-1\},\quad S_{r}=\{M-m_{\Delta}+1,M-m_{\Delta}+2,\ldots,M\}.$ (5)
Let $T_{\Delta}$ be a positive interger,
$L_{\Delta}=2\times T_{\Delta}\times m_{\Delta},\quad h=\frac{2\pi}{L_{\Delta}},\quad x_{j}=(j-1)h,\quad j=1,2,\ldots,L_{\Delta}.$ (6)
Then let
$J_{\Delta,1}=\{1,2,\ldots,m_{\Delta}\},~{}J_{\Delta,2}=\{\frac{L_{\Delta}}{2}+1,\frac{L_{\Delta}}{2}+2,\ldots,\frac{L_{\Delta}}{2}+m_{\Delta}\},~{}J_{\Delta}=J_{\Delta,1}\cup J_{\Delta,2}.$ (7)
For $j\in J_{\Delta}$, take
$g(x_{j})=\left\{\begin{aligned} &f(t_{M-m_{\Delta}+j}),&j\in J_{\Delta,1},\\ &f(t_{-M+j-\frac{L}{2}-1}),&j\in J_{\Delta,2}.\end{aligned}\right.$ (8)
Then we calculate $g_{c}(x_{j}),j=1,2,\ldots,L_{\Delta}$ according to the method in section 3. Finally, we will obtain a function $f_{c}$ with a period of $2+\lambda$, where
$\lambda=\frac{L_{\Delta}-2m_{\Delta}+2}{2M},$ (9)
and the data of the extension function $f_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\frac{L_{\Delta}}{2}-m_{\Delta}$ in one period $[-1,1+\lambda)$ is
$f_{c}(t_{\ell})=\left\{\begin{aligned} &f(t_{\ell}),&|\ell|\leq M,\\ &g_{c}(x_{m_{\Delta}+\ell-M+1}),&\ell>M.\end{aligned}\right.$ (10)
Fig. 2 shows the basic process of the boundary interval algorithm using the function $f(x)=x^{2}$ as an example. Here, in order to make the graph look clearer, we use a relatively small number of nodes. Figure 1 (a) shows the given data (indicated by the black dotted line) and the boundary data selected for the calculation extension (indicated by the red circle). Figure 1 (b) shows the $g_{c}(x_{j})$ calculated from the boundary data. The data represented by the blue circle is the subsequent extension data required. Figure 1 (c) shows the data of $f_{c}$ in one period formed by combining the original data with the calculated extension data. It can be seen from the calculation process that the calculation of the extension part only depends on the boundary interval data.
(a) Given data and selected boundary data
(b) $g_{c}(x_{j})$ with $T_{\Delta}=4$
(c) $f_{c}(t_{\ell})$
Figure 2: Calculation process of a periodic extension of $f(x)=x^{2}$.
After obtaining the data in one period, the subsequent algorithm can be implemented through FFT. Therefore, whether the algorithm can achieve satisfactory efficiency depends on the computational complexity of the extension part, that is, the size of the parameters $T_{\Delta}$ and $m_{\Delta}$. We will give their selection schemes through testing in Section 4.
3 The calculation of $g_{c}$
Including appropriate normalisation, we let
$\phi_{k}(x)=\frac{1}{\sqrt{L_{\Delta}}}e^{\text{i}kx},$ (11)
Assume $g(x_{j})$ is given by (8), then the Fourier extension $g_{c}(x)$ is defined to be
$g_{c}(x)=\sum_{k=-n_{\Delta}}^{n_{\Delta}}{\bf c}_{k}\phi_{k}(x).$ (12)
where the coefficients ${\bf c}_{k}$ are determined by
$\min_{({\bf c}_{k})}\sum_{j\in J_{\Delta}}\left|\sum_{k=-n_{\Delta}}^{n_{\Delta}}{\bf c}_{k}\phi(x_{j})-g(x_{j})\right|^{2}.$ (13)
If we let
[…]
$l=1,2,\ldots,2m_{\Delta},k=-n_{\Delta},\ldots,n_{\Delta}.$
Then the solution of (13) can be obtained by solving the rectangular system
$A{\bf c}\approx b$ (15)
in a least squares sense. If we let
$N_{\Delta}=\min\{2m_{\Delta},2n_{\Delta}+1\},$ (16)
then the SVD of $A$ can be given as
$A=\sum_{i=1}^{N_{\Delta}}{\bf u}_{i}\sigma_{i}{\bf v}^{T}_{i},$ (17)
where $\sigma_{i}$ are the singular values of $A$ such that
$\sigma_{1}\geq\ldots\geq\sigma_{N_{\Delta}}\geq 0,$ (18)
while ${\bf u}_{i}$ and ${\bf v}_{i}$ are the left and right singular vectors of $A$, respectively. We define
$\gamma=\frac{m_{\Delta}}{n_{\Delta}}.$ (19)
Fig. 3 shows the singular values of $A$ calculated by SVD techniques for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$ in MATLAB. It can be seen that the first approximately $\mathcal{O}(n_{\Delta}/T_{\Delta})$ are close to $1$ and then the singular values rapidly decay until the true singular values are masked by machine precision errors.
(a) $n_{\Delta}=100$, $T_{\Delta}=6$
(b) $m_{\Delta}=100$, $T_{\Delta}=20$
(c) $m_{\Delta}=100$, $n_{\Delta}=100$
Figure 3: The singular values of $A$ for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$.
The situation of singular values reflects the ill-posedness of (15), regularization needs to be introduced. Referring to the treatment methods in [2], the truncated SVD leas-squares solution to (15) can be given as
${\bf c}_{\tau}=\sum_{\sigma_{i}>\tau}\frac{{\bf u}_{i}^{T}b}{\sigma_{i}}{\bf v}_{i},$ (20)
where the parameter $\tau$ is usually chosen close to the machine precision.
4 Numerical investigation and determination of key parameters
The proposed algorithm involves multiple parameters, and their selection has a substantial impact on the performance of the algorithm. This section will test and analyze them and then provide an optimized scheme.
4.1 Testing the parameter $T_{\Delta}$
Previous research results show that the extension length has a substantial impact on the convergence rate and the resolution power. Generally speaking, a larger extension length can achieve a higher convergence rate but also means a larger resolution constant. For details, refer to [11, 1, 3]. In [1], the authors also point out that the theoretical results may not be witnessed in computations.
For the proposed algorithm, we are more concerned about the influence of $T_{\Delta}$ on the convergence rate, and whether there is any relationship between it and the number of boundary points $m_{\Delta}$, the ratio coefficient $\gamma$, and the frequency of the approximated function. Now we take
$f(t)=\exp(i\pi\omega t)$
to test different parameters. In Fig 4, we give results for the approximation error against $T_{\Delta}$ with different parameters. Here and what follows, the approximation error evaluated as the maximum pointwise error over an equi-sampled grid ten times denser than the one used for construction and the parameter $\tau=1e-14$ (we also tested with $\tau=1e-12,1e-13,1e-15$ and obtained results similar to those shown here.)
(a) $\gamma=1$, $M=500$, $m_{\Delta}=100$.
(b) $\gamma=1$, $M=1000$, $\omega=20$.
(c) $\gamma=2$, $M=500$, $m_{\Delta}=150$.
(d) $\gamma=2$, $M=1000$, $\omega=20$.
(e) $\gamma=1/2$, $M=500$, $m_{\Delta}=50$.
(f) $\gamma=1/2$, $M=1000$, $\omega=20$.
Figure 4: The approximation error against $T_{\Delta}$ for various parameters
Three distinct regions are visible in all figures:
  • A region $I_{1}:=\{T_{\Delta}:1<T_{\Delta}<\hat{T}_{1}\}$, where the error decreases rapidly to near the machine accuracy as $T_{\Delta}$ increases.
  • A region $I_{2}:=\{T_{\Delta}:\hat{T}_{1}<T_{\Delta}<\hat{T}_{2}\}$, where the error is maintained close to the machine accuracy.
  • A region $I_{3}:=\{T_{\Delta}>\hat{T}_{2}\}$, where the error rebounds, and the increase of $T_{\Delta}$ will cause the error to increase instead.
It can be seen that the value of $\hat{T}_{1}$ is influenced by $\gamma$, but not by $\omega$, $m_{\Delta}$, and $M$, while $\hat{T}_{2}$ is influenced by $\gamma$, $M$, and $\omega$. For $\hat{T}_{2}$, we can give an estimate:
$\hat{T}_{2}\sim\frac{M}{\gamma\omega}.$ (21)
For several different values of $\gamma$, we give the approximate values of $\hat{T}_{1}$ (We tested the functions of $\omega$ from $1$ to $50$, and took the mean of the corresponding $T_{\Delta}$ when their error was less than $135e-$ 13 for the first time as the estimate of $\hat{T}_{1}$), which are listed in Table 1.
When the number of nodes $M$ is not large enough, the regions $I_{1}$ and $I_{2}$ are merged into one region. At this time, the error reaches the optimal value near $\hat{T}_{2}$. This phenomenon can be observed from Fig. 5.
Table 1: The approximation value of $\hat{T}_{1}$ for various $\gamma$.
$\gamma$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
$\hat{T}_{1}$ 1.09 1.20 2.3 5.9 12.5 24.6 48.9
(a) $\gamma=1$, $M=100$, $m_{\Delta}=50$.
(b) $\gamma=2$, $M=400$, $m_{\Delta}=100$.
Figure 5: The approximation error against $T_{\Delta}$ for $\hat{T}_{2}<\hat{T}_{1}$.
What will happen if the $T_{\Delta}$ values selected is less than $\hat{T}_{1}$? Figure 6 shows this. We can see that the error will not decrease after reaching a certain threshold, which cannot be improved by increasing the numbers of $M$ and $m_{\Delta}$. It should be noted that this situation also applies to the Fourier extension algorithm using full data in previous literature. In Figure 7, we show the results of the algorithm using full data. It can be seen that when the value of $T$ is smaller than the corresponding $\hat{T}_{1}$, the method cannot achieve machine accuracy. Similarly, this cannot be improved by increasing the number of $M$.
(a) $\omega=20$, $\gamma=1$, $m_{\Delta}=50$.
(b) $\omega=20$, $\gamma=1$, $M=500$.
Figure 6: The approximation error against $T_{\Delta}$ for ${T}_{\Delta}<\hat{T}_{1}$ (If $m_{\Delta}>M$, we take $m_{\Delta}=M$).
(a) $\omega=20$, $\gamma=1$.
(b) $\omega=20$, $\gamma=1/2$.
Figure 7: The approximation error against $M$ for full data FE with $T<\hat{T}_{1}$.
From the above test results, we can see that the selection of $T_{\Delta}$ is crucial to the approximation accuracy. Only when $T_{\Delta}$ is within the interval $[\hat{T}_{1},\hat{T}_{2}]$, the approximation can obtain satisfactory results. The length and position of this interval are affected by the parameter $\gamma$ and the frequency $\omega$ of the approximated function.
4.2 Testing the parameter $m_{\Delta}$
The size of parameter $m_{\Delta}$ determines whether this study is meaningful. This subsection tests the influences of other parameters on it. First, we take $M>{T}_{\Delta}\gamma\omega$ for observation. As can be seen from Fig. 8, as the value of $m_{\Delta}$ increases, the error decreases rapidly and reaches the machine precision after a certain value $\hat{m}_{\Delta}$, while the value of $\hat{m}_{\Delta}$ is independent of $M$ and $\omega$. In Fig. 9, we further show how the error changes with $m_{\Delta}$ for different values of $\gamma$ and $T_{\Delta}$. It can be seen that the value of $\hat{m}_{\Delta}$ decreases with the increase of $T_{\Delta}$. When $\gamma\leq 1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ is very close. When $\gamma>1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ increases rapidly with the increase of $\gamma$. From the perspective of use, we are more concerned about the $\hat{m}_{\Delta}$ value near $\hat{T}_{1}$. We tested this and the results are listed in Table 2.
Table 2: The approximation values of $\hat{m}_{\Delta}$ for various $\gamma$ and $T_{\Delta}$.
$\gamma$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
$T_{\Delta}$ $1.1$ $1.2$ $2.3$ $6$ $12$ $24$ $50$
$\hat{m}_{\Delta}$ 1820 480 58 17 18 18 19
(a) $M=1000$, $\gamma=2$, $T_{\Delta}=6$.
(b) $\omega=20$, $\gamma=2$, $T_{\Delta}=6$.
Figure 8: The approximation error against $m_{\Delta}$ for various $\omega$ and $M$.
(a) $M=2000$, $\gamma=8$, $\omega=20$.
(b) $M=1000$, $\gamma=4$, $\omega=20$.
(c) $M=500$, $\gamma=2$, $\omega=20$.
(d) $M=500$, $\gamma=1$, $\omega=20$.
(e) $M=500$, $\gamma=1/2$, $\omega=20$.
(f) $M=500$, $\gamma=1/4$, $\omega=20$.
Figure 9: The approximation error against $m_{\Delta}$ for various $\gamma$ and $T_{\Delta}$.
(a) $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$
(b) approximation error
Figure 10: The function and its approximation errors against $m_{\Delta}$ for $M=500$, $\gamma=1$, $T_{\Delta}=6$.
It is also worth noting that for functions with internal oscillations and smooth boundaries, an increase in the number of $m_{\Delta}$ will cause the value of $\hat{T}_{2}$ to decrease. At this time, if the value of $M$ is not large enough, the approximation result using a larger $m_{\Delta}$ may be worse. This can be seen from the approximation results for the function $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$ in Figure 10.
4.3 Computational complexity and parameter determination
The computational complexity of this algorithm consists of three parts:
  • Calculation of $c_{\tau}$. The computational effort in this step is mainly concentrated on the singular value decomposition of the matrix $A$, which is related to $m_{\Delta}$ and $\gamma$, which can be estimated as:
    • If $\gamma<1$, the computational complexity is
      $\mathcal{O}\left(m_{\Delta}^{2}n_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma}m_{\Delta}^{3}\right)$ (22)
    • If $\gamma\geq 1$, the computational complexity is
      $\mathcal{O}\left(m_{\Delta}n^{2}_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma^{2}}m_{\Delta}^{3}\right)$ (23)
  • Calculation of $g_{c}$. This step can be achieved by IFFT, the computational complexity is
    $O\left(L_{\Delta}\log(L_{\Delta})\right)$ (24)
  • Calculation of Fourier coefficients of $f_{c}$. It can be achieved by FFT, the computational complexity is
    $O\left(M\log(M)\right).$ (25)
Combining the test results in the previous two subsections, we know that if we want to approximate the function with frequency $\omega$ with machine precision, the parameters $m_{\Delta}$ and $T_{\Delta}$ need to satisfy
$m_{\Delta}\geq\hat{m}_{\Delta},\quad T_{\Delta}\geq\hat{T}_{1}.$ (26)
Accordingly, combined with formula (21), the parameters $L_{\Delta}$ and $M$ need to satisfy
$L_{\Delta}\geq\hat{L}_{\Delta}=:2\times\hat{T}_{1}\times\hat{m}_{\Delta},\quad M\geq\hat{M}=:\hat{T}_{1}\gamma\omega.$ (27)
Referring to the results in Table 1 and Table 2, the value of $\hat{M}$ is very close for different $\gamma$ and the value of $\hat{L}_{\Delta}$ is relatively close when $\gamma=1$ and $2$, but the value of $\hat{L}_{\Delta}$ is much larger when $\gamma$ takes other values. Further combined with the value of $\hat{m}_{\Delta}$, it can be seen that the computational complexity of the method is equivalent when $\gamma=1$ and $\gamma=2$, while the computational complexity increases significantly in other cases. Furthermore, note that a larger $m_{\Delta}$ means a greater probability of encountering high frequencies in the selected boundary interval. Therefore, we recommend using the following parameters
$\gamma=1,T_{\Delta}=6,m_{\Delta}=20,$ (28)
for actual calculations, and parameters
$\gamma=2,T_{\Delta}=2.3,m_{\Delta}=60,$ (29)
can be used as an alternative. Under such parameter settings, the size of the matrix $A$ in equation (15) is small, and we can directly calculate its singular value decomposition. In addition, note that $A$ is fixed in the algorithm, so the SVD of $A$ can be pre-computed and stored.
5 Testing and comparison of algorithm performance
Through the previous tests, we have given specific values for the parameters involved in the algorithm. In this section, we will further test the performance of the algorithm. All the tests are computed on Windows 10 system with Memory 16 GB, CPU Intel(R) Core(TM)i7-8500U CPU1.80GHz by MATLAB 2016b. We will compare the boundary interval algorithm (M1) proposed in this paper with Lyons algorithm (M2) and the Implicit projection algorithm (M3) in [14]. In the other two algorithms, we take the parameters as $T=2,\gamma=2$.
Figure 11 shows execution time for increasing the number of nodes $M$ of three algorithms. According to the parameter settings in (28), the algorithm only needs to perform an SVD of a matrix with a dimension of $40\times 41$, an IFFT of length $L_{\Delta}=2\times 6\times 20=240$, and an FFT of length $12M+$ 1. As $M$ increases, the speed of the present algorithm becomes apparent.
Figure 11: Execution time for increasing the number of nodes $M$, for MATLAB implementations of the Boundary Interval algorithm (M1), the Lyon algorithm (M2) and Implicit projection algorithm (M3).
The accuracy of the solution obtained through three algorithms is shown in Fig. 12 through 14. In order to better show the differences, we divide the function into three cases for comparison:
  • Overall smooth functions:
    $f_{1}(t)=t^{2},\quad f_{2}(t)=\exp(t),\quad f_{3}(t)=e^{\sin(2.7\pi t)+\cos(\pi t)}.$
  • Internal oscillating functions:
    $f_{4}(t)=\frac{1}{1+100t^{2}},\quad f_{5}(t)=\frac{100}{\cos(1+25t^{2})},\quad f_{6}(t)=\text{erf}(100t).$
  • Boundary oscillating functions:
    $f_{7}(t)=\frac{1}{1.1-t^{2}},\quad f_{8}(t)=\text{Ai}(-66-70t),\quad f_{9}(t)=e^{\sin(65.5\pi t-27\pi)-\cos(20.6\pi t)}.$
(a) $f(t)=t^{2}$
(b) approximation error
(c) $f(t)=\exp(t)$
(d) approximation error
(e) $f(t)=\exp(\sin(2.7\pi t)+\cos(\pi t))$
(f) approximation error
Figure 12: Overall smooth functions and their approximation errors against $M$.
As can be seen from Figure 12, for the first case, when the function is overall smooth, the convergence of the present algorithm is slightly better than the other two algorithms, and it reaches the ideal accuracy at a smaller $M$. At this time, the number of nodes $M$ required by the three algorithms to achieve machine accuracy is relatively small, so there is no obvious difference in performance.
But as shown in Figure 13, for the second case, that is, when the function oscillates internally, since the method proposed in this article relies more on boundary data, the impact of this singularity on it is much smaller than the other two algorithms. The number of nodes $M$ required to achieve machine accuracy is significantly reduced (about $1/4$ of the other two algorithms).
For the third case, since the function oscillates at the boundary, the algorithm in this paper cannot avoid this singularity, so we see that the convergence performance of the three algorithms is very close. However, it should be noted that the number of nodes $M$ required by the three algorithms to achieve machine accuracy is very large at this time. Therefore, as shown in Figure 3, this method still has an advantage in computational complexity.
(a) $f(t)=\frac{1}{1+100t^{2}}$
(b) approximation error
(c) $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$
(d) approximation error
(e) $f(t)=\text{erf}(100t)$
(f) approximation error
Figure 13: Internal oscillating functions and their approximation errors against $M$.
(a) $f(t)=\frac{1}{1.1-t^{2}}$
(b) approximation error
(c) $f(t)=\text{Ai}(-66-70t)$
(d) approximation error
(e) $f(t)=\exp(\sin(65.5\pi t-27\pi)-\cos(20.6\pi t))$
(f) approximation error
Figure 14: Boundary oscillating functions and their approximation errors against $M$.
6 Conclusions
A new fast algorithm based on boundary interval data has been proposed for the computation of the Fourier extension approximation to nonperiodic functions. We tested the relevant parameters and provided the setting scheme. The final algorithm has a computational complexity of $O(M\log(M))$ and can reduce the impact of internal singularities on reconstruction. This algorithm can be used as a basic tool for Fourier extension calculation.